The square of a number is 4 more than three times the number.find the number?
Answer by ankor@dixie-net.com(22740)

n = the number
The square of a number is 4 more than three times the number.find the number?
n^2 = 3n + 4
Arrange as a quadratic equation
n^2 - 3n - 4 = 0
factors to:
(n-4)(n+1) =
Two solutions
n = 4
n = -1
Check to see that both solutions satisfy the given statement